### v1

# Invariant Hilbert spaces of distribution vectors in Lie group representations

Abstract.
For every unitary irreducible representation of a nilpotent Lie group we prove that the representation Hilbert space is the only nonzero invariant Hilbert space of distribution vectors.
Key words and phrases:
nilpotent Lie group; smooth vector; distribution vector
2020 Mathematics Subject Classification:
Primary 22E27; Secondary 17B30, 22E25, 46F10
This work was supported by a grant of the Ministry of Research, Innovation and Digitization, CNCS–UEFISCDI, project number PN-IV-P1-PCE-2023-0264, within PNCDI IV
1. Introduction
Hilbert subspaces of topological vector spaces in the sense of Laurent Schwartz [Sch64] provide an abstract version of reproducing kernels of function spaces and they thus play an important role in many areas of analysis and representation theory of Lie groups, as shown for instance in [Ne00]. There has been a recent interest in uniqueness of Hilbert subspaces of spaces of tempered distributions that are invariant to modulations and translations, partly motivated by problems related to time-frequency analysis and modulation spaces, see [ToGuMaRa21], [BaMoPi24], [RaToVi25].
In this paper we prove an abstract uniqueness theorem for Lie group representations, which recovers some of the preceding results in the special case of Schrödinger representations of the Heisenberg groups. Specifically, we show that if $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is an irreducible unitary representation of a connected and simply connected nilpotent Lie group $G$, then ${\mathcal{H}}$ is the only continuously embedded Hilbert space into the space of distribution vectors ${\mathcal{H}}_{-\infty}$ which is nonzero and is invariant to the action of $G$ (Theorem 3.1).
Our approach is completely different from the methods from the earlier papers, inasmuch as we rely on some basic results in representation theory of nilpotent Lie groups. Thus, in Corollary 3.4, we provide a new perspective on [RaToVi25, Thm. 0.3].
2. Preliminaries
The framework of the present paper is provided by the notions of smooth vectors and distribution vectors with respect to unitary representations of Lie groups. Details of this theory can be found in [Ne10], [BeBe10], [BeBe11], [BeBeGa18], and the references therein.
For every locally convex, topological ${\mathbb{C}}$-vector space ${\mathcal{Y}}$, we denote by ${\mathcal{L}}({\mathcal{Y}})$ the associative algebra of continuous linear operators on ${\mathcal{Y}}$. and by ${\mathcal{Y}}^{\prime}$ the space of continuous antilinear functionals on ${\mathcal{Y}}$ endowed with its $w^{*}$-topology, i.e., the topology of pointwise convergence. If $\langle\cdot,\cdot\rangle\colon{\mathcal{Y}}^{\prime}\times{\mathcal{Y}}\to{\mathbb{C}}$ is the corresponding sesquilinear duality pairing (antilinear in its second variable), then for every $T\in{\mathcal{L}}({\mathcal{Y}})$ there is a unique operator $T^{\prime}\in{\mathcal{L}}({\mathcal{Y}}^{\prime})$ satisfying $\langle T^{\prime}\eta,y\rangle=\langle\eta,Ty\rangle$ for all $\eta\in{\mathcal{Y}}^{\prime}$ and $y\in{\mathcal{Y}}$. If ${\mathcal{X}}$ is a Banach space, since the continuous linear operators on ${\mathcal{X}}$ are precisely the bounded operators, we denote ${\mathcal{B}}({\mathcal{X}}):={\mathcal{L}}({\mathcal{X}})$.
Here is our main setting:
  • $G$ is a connected, simply connected nilpotent Lie group
  • $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is an irreducible (continuous) unitary representation;
  • :=${\mathcal{H}}_{\infty}:=\{v\in{\mathcal{H}}\mid\pi(\cdot)v\in\mathcal{C}^{\infty}(G,{\mathcal{H}})\}$ is the Fréchet space of smooth vectors;
  • ${\mathcal{H}}_{-\infty}$ is the space of distribution vectors, i.e., the space of continuous antilinear functionals on ${\mathcal{H}}_{\infty}$, endowed with the $w^{*}$-topology;
  • $\langle\cdot,\cdot\rangle\colon{\mathcal{H}}_{-\infty}\times{\mathcal{H}}_{\infty}\to{\mathbb{C}}$ is the corresponding sesquilinear duality pairing, antilinear in its second variable, which agrees with the scalar product of the Hilbert space ${\mathcal{H}}$;
  • we have the group representations
    $\pi_{\infty}\colon G\to{\mathcal{L}}({\mathcal{H}}_{\infty})\text{ and }\pi_{-\infty}\colon G\to{\mathcal{L}}({\mathcal{H}}_{-\infty})$
    satisfying $\pi_{\infty}(g)=\pi(g)|_{{\mathcal{H}}_{\infty}}$ and $\pi_{-\infty}(g):=(\pi_{\infty}(g^{-1}))^{\prime}$ for every $g\in G$;
  • ${\mathcal{S}}(G)$ is the associative Fréchet algebra of Schwartz functions on $G$ with the convolution product.
Remark 2.1.
In the setting above, the natural continuous injective linear maps
${\mathcal{H}}_{\infty}\hookrightarrow{\mathcal{H}}\hookrightarrow{\mathcal{H}}_{-\infty}$ (2.1)
intertwine the group representations $\pi_{\infty}$, $\pi$, and $\pi_{-\infty}$.
Using integration with respect to the Haar measure of $G$, the unitary representation $\pi$ gives rise to a Banach *-algebra representation $\pi\colon L^{1}(G)\to{\mathcal{B}}({\mathcal{H}})$. Moreover, we have $\pi({\mathcal{S}}(G)){\mathcal{H}}_{\infty}\subseteq{\mathcal{H}}_{\infty}$, and this gives rise to the algebra representation
$\pi_{\infty}\colon{\mathcal{S}}(G)\to{\mathcal{L}}({\mathcal{H}}_{\infty}).$
Also, the map
$\pi_{-\infty}\colon{\mathcal{S}}(G)\to{\mathcal{L}}({\mathcal{H}}_{-\infty}),\quad\pi_{-\infty}(\varphi):=(\pi(\varphi^{*}))^{\prime}$
is an algebra representation.
The maps in (2.1) intertwine the representations $\pi_{\infty}$, $\pi|_{{\mathcal{S}}(G)}$, $\pi_{-\infty}$ of the associative algebra ${\mathcal{S}}(G)$ and we have
$\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{H}}_{-\infty}\subseteq{\mathcal{H}}_{\infty}.$ (2.2)
Lemma 2.2.
If ${\mathcal{V}}\subseteq{\mathcal{H}}_{\infty}$ is a ${\mathbb{C}}$-linear subspace satisfying $\pi_{\infty}({\mathcal{S}}(G)){\mathcal{V}}\subseteq{\mathcal{V}}$ then either ${\mathcal{V}}=\{0\}$ or ${\mathcal{V}}={\mathcal{H}}_{\infty}$, that is, the representation $\pi_{\infty}\colon{\mathcal{S}}(G)\to{\mathcal{L}}({\mathcal{H}}_{\infty})$ is algebraically irreducible.
Proof.
See [Ho77, Cor. 3.4.1], and also [dC87, Cor. 3.4] and [BeBeGa18, Thm. 2.9]. ∎
Our main result Theorem 3.1 can be regarded as a sharpening of the condition of absence of nontrivial closed invariant subspaces for irreducible unitary representations of nilpotent Lie groups. Before continuing our preparations for its proof, we derive the following simple result that has a similar flavour and may hold an independent interest.
Proposition 2.3.
If ${\mathcal{Y}}\subseteq{\mathcal{H}}_{-\infty}$ is a $w^{*}$-closed linear subspace with $\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{Y}}\subseteq{\mathcal{Y}}$, then we have either ${\mathcal{Y}}=\{0\}$ or ${\mathcal{Y}}={\mathcal{H}}_{-\infty}$.
Proof.
Let ${\mathcal{Y}}^{\perp}:=\{v\in{\mathcal{H}}_{\infty}\mid\langle{\mathcal{Y}},v\rangle=\{0\}\}$, which is a closed linear subspace of ${\mathcal{H}}_{\infty}$. Since ${\mathcal{Y}}\subseteq{\mathcal{H}}_{-\infty}$ is a $w^{*}$-closed linear subspace, we obtain a topological linear isomorphism
${\mathcal{Y}}\simeq({\mathcal{H}}_{\infty}/{\mathcal{Y}}^{\perp})^{\prime}.$ (2.3)
On the other hand, the hypothesis $\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{Y}}\subseteq{\mathcal{Y}}$ implies $\pi_{\infty}({\mathcal{S}}(G)){\mathcal{Y}}^{\perp}\subseteq{\mathcal{Y}}^{\perp}$ hence, by Lemma 2.2, we have either ${\mathcal{Y}}^{\perp}=\{0\}$ or ${\mathcal{Y}}^{\perp}={\mathcal{H}}_{\infty}$. Now the assertion follows by (2.3). ∎
The next lemma is a more precise form of [Ho77, Cor. 3.4.2]. We give below its short proof.
Lemma 2.4.
Let $\pi$ be as above and $\sigma\colon G\to{\mathcal{B}}({\mathcal{X}})$ be another unitary representation of $G$. Assume that the linear map $T\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}$ satisfies
$T\pi(\varphi)=\sigma(\varphi)T\quad\text{for all }\varphi\in{\mathcal{S}}(G).$
Then $T$ extends to a bounded operator on ${\mathcal{H}}$ that is a scalar multiple of an isometry.
Proof.
Let $v_{1},v_{2}\in{\mathcal{H}}_{\infty}\setminus\{0\}$ be such that $\langle v_{1},v_{2}\rangle=0$. The orthogonal projection $P:=\langle\cdot,v_{1}\rangle{v_{1}}$ belongs to ${\mathcal{B}}({\mathcal{H}})_{\infty}$ by [BeBe10, Ex. 3.1]. By [Ho77, Thm. 3.4], there is $\varphi\in{\mathcal{S}}(G)$ with $P=\pi(\varphi)$, hence $P=P^{*}=P^{2}=\pi(\psi)$ where $\psi:=\varphi^{*}\ast\varphi$. Then, by hypothesis, $\sigma(\psi)T=TP=\langle\cdot,v_{1}\rangle Tv_{1}$, hence
$\sigma(\psi)Tv_{1}=Tv_{1}\;\;\text{and}\;\;\sigma(\psi)Tv_{2}=0.$
It follows that $Tv_{1}$, respectively $Tv_{2}$, are eigenvectors for the eigenvalues $1$, respectively 00, of the self-adjoint operator $\sigma(\psi)$, and thus they are orthogonal.
Therefore we have obtained that for $v_{1},v_{2}\in{\mathcal{H}}_{\infty}$,
$\langle v_{1},v_{2}\rangle=0\;\Longrightarrow\;\langle Tv_{1},Tv_{2}\rangle=0.$ (2.4)
Let now $v,w\in{\mathcal{H}}_{\infty}$ be arbitrary with $\|v\|=\|w\|=1$. Set $v_{2}=w-\langle w,v\rangle v$. Then $\langle v,v_{2}\rangle=0$ hence, by (2.4), $\langle Tv,Tv_{2}\rangle=0$ , that is,
$\langle Tv,Tw\rangle=\langle v,w\rangle\|Tv\|^{2}.$ (2.5)
Interchanging the roles of $w$ and $v$ we have obtained that $\langle v,w\rangle\|Tv\|^{2}=\langle v,w\rangle\|Tw\|^{2}$. It follows that
$\|Tv\|=\|Tw\|\quad\text{if }\langle v,w\rangle\neq 0.$
If $\langle v,w\rangle=0$ and we put $w_{\epsilon}=(w+\epsilon v)/\|w+\epsilon v\|$ for every $\epsilon>0$, then $\langle v,w_{\epsilon}\rangle\neq 0$. Using the above equality we get that
$\|Tv\|=\|w+\epsilon v\|^{-1}\|T(w+\epsilon v)\|=\|w+\epsilon v\|^{-1}\|Tw+\epsilon Tv\|.$
Letting $\epsilon\to 0$, we get that
$\|Tv\|=\|Tw\|\quad\text{for every}\quad v,w\in{\mathcal{H}}_{\infty},\;\|v\|=\|w\|=1.$
Thus $\lambda_{0}:=\|Tv\|$ is independent on $v\in{\mathcal{H}}_{\infty}$ with $\|v\|=1$. Then by (2.5) we get
$\langle Tv,Tw\rangle=\lambda_{0}^{2}\langle v,w\rangle$ (2.6)
[…]
3. The main result
Theorem 3.1.
Let $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ be an irreducible unitary representation of a connected, simply connected nilpotent Lie group $G$. Let ${\mathcal{X}}\subseteq{\mathcal{H}}_{-\infty}$ be a ${\mathbb{C}}$-linear subspace endowed with the structure of a complex Hilbert space satisfying the following hypotheses:
  1. (H1)
    the inclusion map ${\mathcal{X}}\hookrightarrow{\mathcal{H}}_{-\infty}$ is continuous;
  2. (H2)
    $\pi_{-\infty}(G){\mathcal{X}}\subseteq{\mathcal{X}}$;
  3. (H3)
    for every $g\in G$ the operator $\pi_{\mathcal{X}}(g):=\pi_{-\infty}(g)|_{\mathcal{X}}\colon{\mathcal{X}}\to{\mathcal{X}}$ is an isometry.
Then either ${\mathcal{X}}=\{0\}$ or ${\mathcal{X}}={\mathcal{H}}$.
For the proof of the above theorem we still need the following lemma.
Lemma 3.2.
Let $\Gamma$ be a topological group, ${\mathcal{X}}$ a complex Hilbert space, $\rho\colon\Gamma\to{\mathcal{B}}({\mathcal{X}})$ a mapping, and for every $x\in{\mathcal{X}}$ denote $\rho^{x}\colon\Gamma\to{\mathcal{X}}$, $\rho^{x}(g):=\rho(g)x$. Assume that the following conditions are satisfied:
  1. (a)
    For every $g,h\in\Gamma$ we have $\rho(gh)=\rho(g)\rho(h)$ and moreover $\rho({\bf 1})={\rm id}_{\mathcal{X}}$.
  2. (b)
    For every $g\in\Gamma$ the operator $\rho(g)\colon{\mathcal{X}}\to{\mathcal{X}}$ is an isometry.
  3. (c)
    There exists a subset $A\subseteq{\mathcal{X}}$ that spans a dense linear subspace of ${\mathcal{X}}$ and $\rho^{x}\colon\Gamma\to{\mathcal{X}}$ is continuous for every $x\in A$.
Then $\rho$ is a continuous unitary representation of the topological group $\Gamma$.
Proof.
For every $g\in\Gamma$ we have ${\rm id}_{\mathcal{X}}=\rho({\bf 1})=\rho(g)\rho(g^{-1})$, hence $\rho(g)\in{\mathcal{B}}({\mathcal{X}})$ is a surjective isometry. that is, a unitary operator. It remains to prove that $\rho^{x}\colon\Gamma\to{\mathcal{X}}$ is continuous for every $x\in{\mathcal{X}}$. (See e.g., [Ne10, Lemma 5.2].) To this end we consider the set ${\mathcal{X}}_{0}:=\{x\in{\mathcal{X}}\mid\rho^{x}\in{\mathcal{C}}(\Gamma,{\mathcal{X}})\}$. It is clear that ${\mathcal{X}}_{0}$ is a linear subspace of ${\mathcal{X}}$ and $A\subseteq{\mathcal{X}}_{0}$, hence ${\mathcal{X}}_{0}$ is dense in ${\mathcal{X}}$. On the other hand, for arbitrary $x,y\in{\mathcal{X}}$ and $g\in\Gamma$ we have $\|\rho^{x}(g)-\rho^{y}(g)\|=\|\rho(g)x-\rho(g)y\|=\|x-y\|$. This directly implies that ${\mathcal{X}}_{0}$ is closed in ${\mathcal{X}}$, hence ${\mathcal{X}}_{0}={\mathcal{X}}$. ∎
Proof of Theorem 3.1.
We assume ${\mathcal{X}}\neq\{0\}$. For every $v\in{\mathcal{H}}_{\infty}$ we have $\overline{\langle\cdot,v\rangle}\in({\mathcal{H}}_{-\infty})^{\prime}$, hence $\overline{\langle\cdot,v\rangle}|_{\mathcal{X}}\in{\mathcal{X}}^{\prime}$ by (H1). Since ${\mathcal{X}}$ is a complex Hilbert space, there exists a uniquely determined vector $\iota(v)\in{\mathcal{X}}$ satisfying
$\langle\cdot,v\rangle|_{\mathcal{X}}=\langle\cdot,\iota(v)\rangle_{\mathcal{X}},$ (3.1)
where $\langle\cdot,\cdot\rangle_{\mathcal{X}}$ denotes the scalar product of ${\mathcal{X}}$. We thus obtain a linear operator
$\iota\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}.$
The equality (3.1) shows that $\iota$ is the adjoint of the inclusion map ${\mathcal{X}}\hookrightarrow{\mathcal{H}}_{-\infty}$. The range of the operator $\iota$ is dense in ${\mathcal{X}}$ since the above inclusion map is injective. In fact, if $x_{0}\in{\mathcal{X}}$ satisfies $x_{0}\perp\iota({\mathcal{H}}_{\infty})$ in ${\mathcal{X}}$ then, by (3.1), we have $\langle x_{0},v\rangle=0$ for every $v\in{\mathcal{H}}_{\infty}$, that is, $x_{0}=0$.
In order to prove that $\iota$ is a continuous operator, one can use the Closed Graph Theorem since ${\mathcal{H}}_{\infty}$ is a Fréchet space and ${\mathcal{X}}$ is a Hilbert space. Specifically, if $\lim\limits_{n\to\infty}v_{n}=v$ in ${\mathcal{H}}_{\infty}$ and $\lim\limits_{n\to\infty}\iota(v_{n})=x$ in ${\mathcal{X}}$, then $\iota(v)=x$ by (3.1).
For every $g\in G$, $v\in{\mathcal{H}}_{\infty}$, and $x\in{\mathcal{X}}$ we have
$\langle x,\iota(\pi_{\infty}(g)v)\rangle_{\mathcal{X}}$ $=\langle x,\pi_{\infty}(g)v\rangle=\langle\pi_{-\infty}(g^{-1})x,v\rangle=\langle\pi_{\mathcal{X}}(g^{-1})x,v\rangle$
$=\langle\pi_{\mathcal{X}}(g^{-1})x,\iota(v)\rangle_{\mathcal{X}}=\langle x,\pi_{\mathcal{X}}(g)\iota(v)\rangle_{\mathcal{X}}$
hence
$(\forall g\in G,v\in{\mathcal{H}}_{\infty})\quad\iota(\pi_{\infty}(g)v)=\pi_{\mathcal{X}}(g)\iota(v).$ (3.2)
Since $\iota\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}$ is a continuous linear operator, and the mapping $G\ni g\mapsto\pi_{\infty}(g)v\in{\mathcal{H}}_{\infty}$ is smooth as $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is a continuous representation (cf. [Po72, Prop. 1.2]), it also follows by (3.2) that for every $x\in\iota({\mathcal{H}}_{\infty})\subseteq{\mathcal{X}}$ the mapping $\rho_{\mathcal{X}}^{x}\colon G\to{\mathcal{X}}$, $\pi_{\mathcal{X}}^{x}(g):=\pi_{\mathcal{X}}(g)x$, is smooth, and in particular continuous. We proved above that $\iota({\mathcal{H}}_{\infty})$ is dense in ${\mathcal{X}}$, hence by, (H3), Lemma 3.2 is applicable and ensures that $\pi_{\mathcal{X}}\colon G\to{\mathcal{B}}({\mathcal{X}})$ is a continuous unitary representation, and moreover $\iota({\mathcal{H}}_{\infty})\subseteq{\mathcal{X}}_{\infty}$. Then, using (3.2) once again, we obtain
$(\forall\varphi\in{\mathcal{S}}(G))\quad\iota\circ\pi_{\infty}(\varphi)=(\pi_{\mathcal{X}})_{\infty}(\varphi)\circ\iota\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}_{\infty}.$
This implies by Lemma 2.4 that there exist $\lambda_{0}\in{\mathbb{R}}$ and an isometry $\iota_{0}\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}$ with $\iota=\lambda_{0}\iota_{0}$. Since $\iota$ is has dense range in ${\mathcal{X}}$ and thus is non-trivial when ${\mathcal{X}}\neq\{0\}$, $\lambda_{0}\neq 0$ and we may assume that $\lambda_{0}>0$, by replacing $\iota_{0}$ by $-\iota_{0}$ if necessary.
Using again that the range of $\iota$ is dense in ${\mathcal{X}}$ it follows that $\iota_{0}\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}$ extends by continuity to a unitary operator $\iota_{0}\colon{\mathcal{H}}\to{\mathcal{X}}$ that intertwines the representations $\pi$ and $\pi_{\mathcal{X}}$. In particular, the unitary representation $\pi_{\mathcal{X}}$ is irreducible.
On the other hand, we recall from (2.2) that $\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{H}}_{-\infty}\subseteq{\mathcal{H}}_{\infty}$, in particular $\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{X}}\subseteq{\mathcal{H}}_{\infty}$, while the hypothesis (H2) along with the continuity of the representation $\pi_{\mathcal{X}}(\cdot)=\pi_{-\infty}(\cdot)|_{\mathcal{X}}$ implies by (2.2) again that $\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{X}}\subseteq{\mathcal{X}}_{\infty}$. Thus
$\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{X}}\subseteq{\mathcal{X}}_{\infty}\cap{\mathcal{H}}_{\infty}$
and then the property ${\mathcal{X}}\neq\{0\}$ shows that ${\mathcal{X}}_{\infty}\cap{\mathcal{H}}_{\infty}\neq\{0\}$. As above, by the hypothesis (H2) again we also have $\pi_{\infty}({\mathcal{S}}(G))({\mathcal{X}}_{\infty}\cap{\mathcal{H}}_{\infty})\subseteq{\mathcal{X}}_{\infty}\cap{\mathcal{H}}_{\infty}$. Then, by Lemma 2.2 applied for both irreducible representations $\pi$ and $\pi_{\mathcal{X}}$, we obtain ${\mathcal{H}}_{\infty}={\mathcal{X}}_{\infty}\cap{\mathcal{H}}_{\infty}={\mathcal{X}}_{\infty}$, that is, ${\mathcal{H}}_{\infty}={\mathcal{X}}_{\infty}$.
By another application of Lemma 2.4, the identity map ${\rm id}_{{\mathcal{H}}_{\infty}}\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}$ extends to ${\mathcal{H}}$ and its extension is a scalar multiple of an isometry. Since ${\mathcal{H}}_{\infty}$ is dense in ${\mathcal{H}}$, and ${\mathcal{X}}_{\infty}$ is dense in ${\mathcal{X}}$, we get that ${\mathcal{H}}\subset{\mathcal{X}}$, this extension is a multiple of an unitary operator and it is nothing else than ${\rm id}_{{\mathcal{H}}}\colon{\mathcal{H}}\to{\mathcal{X}}$. That is, ${\mathcal{X}}={\mathcal{H}}$ as sets, and the scalar products of these Hilbert spaces are a suitable scalar multiple of each other. ∎
Remark 3.3.
In the setting of Theorem 3.1, if we assume (H2), then (H3) is equivalent to the following, apparently weaker, condition:
  1. (H4)
    there exist a constant $C>0$ and an equivalent Hilbert space norm $\|\cdot\|^{\prime}_{\mathcal{X}}$ such that for every $g\in G$ and $x\in{\mathcal{X}}$ we have $\|\pi_{\mathcal{X}}(g)x\|^{\prime}_{\mathcal{X}}\leq C\|x\|^{\prime}_{\mathcal{X}}$.
In fact, clearly (H3$\Rightarrow$(H4) with $\|\cdot\|^{\prime}_{\mathcal{X}}=\|\cdot\|^{\prime}_{\mathcal{X}}$ and $C=1$. For the converse implication, let us assume (H4) and denote by $G_{d}$ the group $G$ endowed with its discrete topology. Since $G_{d}$ is a nilpotent group, it is solvable hence amenable. See e.g., [Gr69, page 9]. Then, by [Gr69, Thm. 3.4.1] and its proof, one can define an equivalent Hilbert space norm $\|\cdot\|_{\mathcal{X}}$ on ${\mathcal{X}}$ for which the representation $\pi_{\mathcal{X}}\colon G_{d}\to{\mathcal{B}}({\mathcal{X}})$ is unitary, that is, (H3) holds true.
In the special case when $G$ is a Heisenberg group and $\pi$ is a Schrödinger representation as in the proof of Corollary 3.4 below, the above remark provides a new approach to the result of [RaToVi25, Lemma 1.3].
We are now in a position to provide a new proof of [RaToVi25, Thm. 0.3] as follows.
Corollary 3.4.
Let ${\mathcal{X}}\hookrightarrow{\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$ be a continuous inclusion of a nonzero complex Hilbert space into the space of tempered distributions on ${\mathbb{R}}^{d}$, satisfying the following condition: there exists a constant $C_{0}>0$ such that for all $f\in{\mathcal{X}}$ and $x,\xi\in{\mathbb{R}}^{d}$ we have $f(\cdot-x){\rm e}^{{\rm i}\langle\cdot,\xi\rangle}\in{\mathcal{X}}$ and $\|f(\cdot-x){\rm e}^{{\rm i}\langle\cdot,\xi\rangle}\|_{\mathcal{X}}\leq C_{0}\|f\|_{\mathcal{X}}$.
Then ${\mathcal{X}}=L^{2}({\mathbb{R}}^{d})$ as vector spaces, and the norms $\|\cdot\|_{\mathcal{X}}$ and $\|\cdot\|_{L^{2}({\mathbb{R}}^{d})}$ are equivalent.
Proof.
We consider the Heisenberg group $G=H_{d}={\mathbb{R}}^{d}\times{\mathbb{R}}^{d}\times{\mathbb{R}}$ with its group operation
$(x,\xi,t)\cdot(y,\eta,s):=(x+y,\xi+\eta,t+s+(\langle\xi,y\rangle-\langle\eta,x\rangle)/2)$
and the Schrödinger representation $\pi\colon H_{d}\to{\mathcal{B}}(L^{2}({\mathbb{R}}^{d}))$ given by
$(\pi(x,\xi,t)f)(y)={\rm e}^{{\rm i}(\langle\xi,y\rangle+t+\langle\xi,x\rangle/2)}f(y+x).$
(See e.g., [CG90, Ex. 2.2.6].) Then we have ${\mathcal{H}}_{\infty}={\mathcal{S}}({\mathbb{R}}^{d})$ hence ${\mathcal{H}}_{-\infty}={\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$. Thus the assumption on ${\mathcal{X}}$ implies that the hypotheses (H1) and (H2) in Theorem 3.1, as well as (H4) in Remark 3.3 are satisfied. Now the assertion follows by Theorem 3.1 along with Remark 3.3. ∎
Acknowledgment
We wish to thank Dr. Shubham Bais for kindly providing an electronic copy of [BaMoPi24].