### v2

# Invariant Hilbert spaces of distribution vectors in Lie group representations

Abstract.
For every unitary irreducible representation of a Lie group we prove that the representation Hilbert space is the only nonzero invariant Hilbert space of distribution vectors.
Key words and phrases:
nilpotent Lie group; smooth vector; reproducing kernel
2020 Mathematics Subject Classification:
Primary 22E27; Secondary 17B30, 22E25, 46F10
This work was supported by a grant of the Ministry of Research, Innovation and Digitization, CNCS–UEFISCDI, project number PN-IV-P1-PCE-2023-0264, within PNCDI IV
1. Introduction
Hilbert subspaces of topological vector spaces in the sense of Laurent Schwartz [Sch64] provide an abstract version of reproducing kernels of function spaces and they thus play an important role in many areas of mathematical physics, analysis, and representation theory of Lie groups, as shown for instance in [Ne00]. There has been a recent interest in uniqueness of Hilbert subspaces of spaces of tempered distributions that are invariant to modulations and translations, partly motivated by problems related to time-frequency analysis and modulation spaces, see [ToGuMaRa21], [BaMoPi24], [RaToVi25].
In this paper we prove an abstract uniqueness theorem for Lie group representations, which recovers some of the preceding results in the special case of Schrödinger representations of the Heisenberg groups. Specifically, we show that if $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is an irreducible unitary representation of a Lie group $G$, then $\mathcal{H}$} is the only continuously embedded Hilbert space into the space of distribution vectors $\mathcal{H}$$^{-\infty}$ which is nonzero and is invariant to the action of $G$ (Theorem 3.1). In section 4 we give a version of that result replacing the action of the group by the action of the Schwartz algebra on a nilpotent Lie group.
As explained in [ToGuMaRa21], the minimality properties of the Feichtinger algebra provided the initial motivation for uniqueness results on Hilbert spaces that are invariant to translations and modulations. In that spirit, one used the realization of the Hilbert space as $L^{2}({\mathbb{R}}^{d})$, which facilitated the techniques of distribution theory and Hermite functions. Our approach is completely different, inasmuch as we rather rely on some basic results on representations of Lie groups and thus, in Corollary 3.4, we provide a new perspective on [RaToVi25, Thm. 0.3].
2. Preliminaries
The framework of the present paper is provided by the notions of smooth vectors and distribution vectors with respect to unitary representations of Lie groups. Details of this theory can be found in [Ne10], [BeBe10], [BeBe11], [BeBeGa18], and the references therein.
For every locally convex, topological $\mathbb{C}$}-vector space $\mathcal{Y}$}, we denote by ${\mathcal{L}}({\mathcal{Y}})$ the associative algebra of continuous linear operators on $\mathcal{Y}$}. and by $\mathcal{Y}$$^{\prime}$ the space of continuous antilinear functionals on $\mathcal{Y}$} endowed with its $w^{*}$-topology, i.e., the topology of pointwise convergence. If $\langle\cdot,\cdot\rangle\colon{\mathcal{Y}}^{\prime}\times{\mathcal{Y}}\to{\mathbb{C}}$ is the corresponding sesquilinear duality pairing (antilinear in its second variable), then for every $T\in{\mathcal{L}}({\mathcal{Y}})$ there is a unique operator $T^{\prime}\in{\mathcal{L}}({\mathcal{Y}}^{\prime})$ satisfying $\langle T^{\prime}\eta,y\rangle=\langle\eta,Ty\rangle$ for all $\eta\in{\mathcal{Y}}^{\prime}$ and $y\in{\mathcal{Y}}$. If $\mathcal{X}$} is a Banach space, since the continuous linear operators on $\mathcal{X}$} are precisely the bounded operators, we denote ${\mathcal{B}}({\mathcal{X}}):={\mathcal{L}}({\mathcal{X}})$.
We need a version of Schur’s lemma for unbounded intertwining operators.
Definition 2.1.
Let $G$ be a locally compact group and $\pi\colon G\to{\mathcal{B}}({\mathcal{H}}_{j})$, $j=1,2$, two unitary representations of the group $G$. An unbounded operator $S$ from ${\mathcal{H}}_{1}$ to ${\mathcal{H}}_{2}$ with domain $D(S)\subseteq{\mathcal{H}}_{1}$ is an intertwining operator for the representations $\pi_{1}$ and $\pi_{2}$ if for every $g\in G$ and $x\in D(S)$ we have $\pi(g)x\in D(S)$ and
$S\pi_{1}(g)x=\pi_{2}(g)Sx.$
Lemma 2.2.
Let $G$ be a locally compact group and $\pi\colon G\to{\mathcal{B}}({\mathcal{H}}_{j})$, $j=1,2$, two unitary representations of the group $G$ with $\pi_{1}$ irreducible. Assume that $S$ is non-zero closable intertwining operator for $\pi_{1}$ and $\pi_{2}$. Then the closure $\overline{S}$ of $S$ is a bounded operator from ${\mathcal{H}}_{1}$ to ${\mathcal{H}}_{2}$ and there is $\lambda>0$ such that $\lambda\overline{S}\colon{\mathcal{H}}_{1}\to{\mathcal{H}}_{2}$ is an isometry.
Proof.
This is a direct consequence of [Ne00, Lemma II.2.6 and Thm. II.2.8]. ∎
In the present paper, unless otherwise mentioned, we use the following notation.
  • $G$ is a Lie group;
  • $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is a (continuous) unitary representation;
  • ${\mathcal{H}}^{\infty}:=\{v\in{\mathcal{H}}\mid\pi(\cdot)v\in\mathcal{C}^{\infty}(G,{\mathcal{H}})\}$ is the Fréchet space of smooth vectors;
  • $\mathcal{H}$$^{-\infty}$ is the space of distribution vectors, i.e., the space of continuous antilinear functionals on $\mathcal{H}$$^{\infty}$, endowed with the $w^{*}$-topology;
  • $\langle\cdot,\cdot\rangle\colon{\mathcal{H}}^{-\infty}\times$$\mathcal{H}$$^{\infty}\to{\mathbb{C}}$ is the corresponding sesquilinear duality pairing, antilinear in its second variable, which agrees with the scalar product of the Hilbert space $\mathcal{H}$};
  • we have the group representations
[…]
    satisfying $\pi^{\infty}(g)=\pi(g)|_{{\mathcal{H}}^{\infty}}$ and $\pi^{-\infty}(g):=(\pi^{\infty}(g^{-1}))^{\prime}$ for every $g\in G$;
  • $C_{0}^{\infty}(G)$ is the space of compactly supported smooth functions on $G$.
Remark 2.3.
In the setting above, the natural continuous injective linear maps
$\mathcal{H}$$^{\infty}\hookrightarrow{\mathcal{H}}\hookrightarrow$$\mathcal{H}$$^{-\infty}$ (2.1)
intertwine the group representations $\pi^{\infty}$, $\pi$, and $\pi^{-\infty}$.
Using integration with respect to a Haar measure of $G$, the unitary representation $\pi$ gives rise to a Banach *-algebra representation $\pi\colon L^{1}(G)\to{\mathcal{B}}({\mathcal{H}})$. Moreover, we have $\pi(C_{0}^{\infty}(G)){\mathcal{H}}^{\infty}\subseteq{\mathcal{H}}^{\infty}$ with dual map
$\pi^{-\infty}\colon C_{0}^{\infty}(G)\to{\mathcal{L}}({\mathcal{H}}^{-\infty}),\quad\pi^{-\infty}(\varphi):=(\pi(\varphi^{*}))^{\prime}.$
Then
$\pi^{-\infty}(C_{0}^{\infty}(G)){\mathcal{H}}^{-\infty}={\mathcal{H}}^{\infty}.$ (2.2)
by [DM78, Thm. 3.1 and 3.2].
3. The main result
Theorem 3.1.
Let $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ be an irreducible unitary representation of a Lie group $G$. Let $\mathcal{X}$$\subseteq{\mathcal{H}}^{-\infty}$ be a $\mathbb{C}$}-linear subspace endowed with the structure of a complex Hilbert space satisfying the following hypotheses:
  1. (H1)
    the inclusion map $\mathcal{X}$$\hookrightarrow{\mathcal{H}}^{-\infty}$ is continuous;
  2. (H2)
    $\pi^{-\infty}(G){\mathcal{X}}\subseteq{\mathcal{X}}$;
  3. (H3)
    for every $g\in G$ the operator $\pi_{\mathcal{X}}(g):=\pi^{-\infty}(g)|_{\mathcal{X}}\colon{\mathcal{X}}\to{\mathcal{X}}$ is an isometry.
Then either ${\mathcal{X}}=\{0\}$ or $\mathcal{X}$$\mathcal{H}$}.
For the proof of the above theorem we need the following lemma.
Lemma 3.2.
Let $\Gamma$ be a topological group, $\mathcal{X}$} a complex Hilbert space, $\rho\colon\Gamma\to{\mathcal{B}}({\mathcal{X}})$ a mapping, and for every $x\in{\mathcal{X}}$ denote $\rho^{x}\colon\Gamma\to{\mathcal{X}}$, $\rho^{x}(g):=\rho(g)x$. Assume that the following conditions are satisfied:
  1. (a)
    For every $g,h\in\Gamma$ we have $\rho(gh)=\rho(g)\rho(h)$ and moreover $\rho({\bf 1})={\rm id}_{\mathcal{X}}$.
  2. (b)
    For every $g\in\Gamma$ the operator $\rho(g)\colon{\mathcal{X}}\to{\mathcal{X}}$ is an isometry.
  3. (c)
    There exists a subset $A\subseteq{\mathcal{X}}$ that spans a dense linear subspace of $\mathcal{X}$} and $\rho^{x}\colon\Gamma\to{\mathcal{X}}$ is continuous for every $x\in A$.
Then $\rho$ is a continuous unitary representation of the topological group $\Gamma$.
Proof.
For every $g\in\Gamma$ we have ${\rm id}_{\mathcal{X}}=\rho({\bf 1})=\rho(g)\rho(g^{-1})$, hence $\rho(g)\in{\mathcal{B}}({\mathcal{X}})$ is a surjective isometry. that is, a unitary operator. It remains to prove that $\rho^{x}\colon\Gamma\to{\mathcal{X}}$ is continuous for every $x\in{\mathcal{X}}$. (See e.g., [Ne10, Lemma 5.2].) To this end we consider the set ${\mathcal{X}}_{0}:=\{x\in{\mathcal{X}}\mid\rho^{x}\in{\mathcal{C}}(\Gamma,{\mathcal{X}})\}$. It is clear that ${\mathcal{X}}_{0}$ is a linear subspace of $\mathcal{X}$} and $A\subseteq{\mathcal{X}}_{0}$, hence ${\mathcal{X}}_{0}$ is dense in $\mathcal{X}$}. On the other hand, for arbitrary $x,y\in{\mathcal{X}}$ and $g\in\Gamma$ we have $\|\rho^{x}(g)-\rho^{y}(g)\|=\|\rho(g)x-\rho(g)y\|=\|x-y\|$. This directly implies that ${\mathcal{X}}_{0}$ is closed in $\mathcal{X}$}, hence ${\mathcal{X}}_{0}={\mathcal{X}}$. ∎
Proof of Theorem 3.1.
We assume ${\mathcal{X}}\neq\{0\}$. For every $v\in{\mathcal{H}}^{\infty}$ we have $\overline{\langle\cdot,v\rangle}\in({\mathcal{H}}^{-\infty})^{\prime}$, hence $\overline{\langle\cdot,v\rangle}|_{\mathcal{X}}\in{\mathcal{X}}^{\prime}$ by (H1). Since $\mathcal{X}$} is a complex Hilbert space, there exists a uniquely determined vector $\iota(v)\in{\mathcal{X}}$ satisfying
$\langle\cdot,v\rangle|_{\mathcal{X}}=\langle\cdot,\iota(v)\rangle_{\mathcal{X}},$ (3.1)
where $\langle\cdot,\cdot\rangle_{\mathcal{X}}$ denotes the scalar product of $\mathcal{X}$}. We thus obtain a linear operator
$\iota\colon{\mathcal{H}}^{\infty}\to{\mathcal{X}},$
such that
$\iota(v)=v\quad\text{for }v\in{\mathcal{H}}^{\infty}\cap{\mathcal{X}}.$
The equality (3.1) shows that $\iota$ is the adjoint of the inclusion map $\mathcal{X}$$\hookrightarrow{\mathcal{H}}^{-\infty}$. The range of the operator $\iota$ is dense in $\mathcal{X}$} since the above inclusion map is injective. In fact, if $x_{0}\in{\mathcal{X}}$ satisfies $x_{0}\perp\iota({\mathcal{H}}^{\infty})$ in $\mathcal{X}$} then, by (3.1), we have $\langle x_{0},v\rangle=0$ for every $v\in{\mathcal{H}}^{\infty}$, that is, $x_{0}=0$.
In order to prove that $\iota$ is a continuous operator, one can use the closed graph theorem since $\mathcal{H}$$^{\infty}$ is a Fréchet space and $\mathcal{X}$} is a Hilbert space. Specifically, if $\lim\limits_{n\to\infty}v_{n}=v$ in $\mathcal{H}$$^{\infty}$ and $\lim\limits_{n\to\infty}\iota(v_{n})=x$ in $\mathcal{X}$}, then $\iota(v)=x$ by (3.1).
For every $g\in G$, $v\in{\mathcal{H}}^{\infty}$ and $x\in{\mathcal{X}}$ we have
$\langle x,\iota(\pi^{\infty}(g)v)\rangle_{\mathcal{X}}$ $=\langle x,\pi^{\infty}(g)v\rangle=\langle\pi^{-\infty}(g^{-1})x,v\rangle=\langle\pi_{\mathcal{X}}(g^{-1})x,v\rangle$
$=\langle\pi_{\mathcal{X}}(g^{-1})x,\iota(v)\rangle_{\mathcal{X}}=\langle x,\pi_{\mathcal{X}}(g)\iota(v)\rangle_{\mathcal{X}},$
hence
$(\forall g\in G,v\in{\mathcal{H}}^{\infty})\quad\iota(\pi^{\infty}(g)v)=\pi_{\mathcal{X}}(g)\iota(v).$ (3.2)
Since $\iota\colon{\mathcal{H}}^{\infty}\to{\mathcal{X}}$ is a continuous linear operator and the mapping $G\ni g\mapsto\pi^{\infty}(g)v\in{\mathcal{H}}^{\infty}$ is smooth, as $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is a continuous representation (cf. [Po72, Prop. 1.2]), it also follows by (3.2) that, for every $x\in\iota({\mathcal{H}}^{\infty})\subseteq{\mathcal{X}}$, the mapping $\rho_{\mathcal{X}}^{x}\colon G\to{\mathcal{X}}$, $\pi_{\mathcal{X}}^{x}(g):=\pi_{\mathcal{X}}(g)x$, is smooth and in particular continuous. We proved above that $\iota({\mathcal{H}}^{\infty})$ is dense in $\mathcal{X}$}, hence by (H3), Lemma 3.2 is applicable and ensures that $\pi_{\mathcal{X}}\colon G\to{\mathcal{B}}({\mathcal{X}})$ is a continuous unitary representation. Moreover, (3.2) shows that $\iota({\mathcal{H}}^{\infty})\subseteq{\mathcal{X}}^{\infty}$, hence
${\mathcal{H}}^{\infty}\cap{\mathcal{X}}\subseteq{\mathcal{X}}^{\infty}.$ (3.3)
On the other hand, we recall from (2.2) that $\pi^{-\infty}(C_{0}^{\infty}(G))){\mathcal{H}}^{-\infty}={\mathcal{H}}^{\infty}$; in particular, $\pi_{{\mathcal{X}}}(C_{0}^{\infty}(G)){\mathcal{X}}\subseteq{\mathcal{H}}^{\infty}$. Then Dixmier-Malliavin’s theorem [DM78, Thm. 3.2] for the continuous representation $\pi_{\mathcal{X}}$ shows that $\mathcal{X}$$^{\infty}\subseteq{\mathcal{H}}^{\infty}$ which, along with (3.3) gives
${\mathcal{X}}^{\infty}={\mathcal{H}}^{\infty}\cap{\mathcal{X}}.$ (3.4)
The subspace $\mathcal{X}$$^{\infty}$ is then contained in $\mathcal{H}$} and dense in $\mathcal{X}$} . Moreover since it is invariant under $\pi(g)$, $g\in G$, and the representation $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is irreducible, it follows that $\mathcal{X}$$^{\infty}$ is dense in $\mathcal{H}$}, as well.
We now regard the operator $\iota$ above as an unbounded operator defined on $\mathcal{H}$} with domain $\mathcal{H}$$^{\infty}$. A vector $x\in{\mathcal{X}}$ belongs to the domain of the adjoint $\iota^{*}$ of $\iota$ if and only if the map $v\mapsto\langle x,\iota(v)\rangle_{\mathcal{X}}=\langle x,v\rangle$ is continuous in the norm of $\mathcal{H}$}. Since the duality pairing $\langle\cdot,\cdot\rangle$ agrees with the scalar product of $\mathcal{H}$}, we obtain that $\mathcal{X}$$^{\infty}$ is contained in the domain of $\iota^{*}$. It follows that $\iota^{*}$ is densely defined, therefore $\iota$ is a closable operator. Using now (3.2) and Lemma 2.2 we get that there are an isometry $\iota_{0}\colon{\mathcal{H}}\to{\mathcal{X}}$ and a constant $\lambda_{0}>0$ such that $\iota=\lambda_{0}\iota_{0}|_{{\mathcal{H}}^{\infty}}$. As the range of $\iota$, hence of $\iota_{0}$, is dense in $\mathcal{X}$}, it follows that the operator $\iota_{0}\colon{\mathcal{H}}\to{\mathcal{X}}$ is unitary. Moreover, since $\iota|_{{\mathcal{X}}^{\infty}}$ is the identity map, and $\mathcal{X}$$^{\infty}$ is dense both in $\mathcal{X}$} and $\mathcal{H}$}, we get that ${\mathcal{H}}={\mathcal{X}}$ as sets and the scalar products of these Hilbert spaces are a suitable scalar multiple of each other. ∎
Remark 3.3.
In the setting of Theorem 3.1, we denote by $G_{d}$ the group $G$ endowed with its discrete topology. If we assume that $G_{d}$ is amenable and (H2) holds true, then (H3) is equivalent to the following, seemingly weaker, condition:
  1. (H4)
    there exist a constant $C>0$ and an equivalent Hilbert space norm $\|\cdot\|^{\prime}_{\mathcal{X}}$ such that for every $g\in G$ and $x\in{\mathcal{X}}$ we have $\|\pi_{\mathcal{X}}(g)x\|^{\prime}_{\mathcal{X}}\leq C\|x\|^{\prime}_{\mathcal{X}}$.
[…]
We note that if the Lie group $G$ is connected, the condition that $G_{d}$ is amenable is equivalent with $G$ being solvable. (See [Pa88, Thm. 3.9].)
In the special case when $G$ is a Heisenberg group and $\pi$ is a Schrödinger representation as in the proof of Corollary 3.4 below, the above remark provides a new approach to the result of [RaToVi25, Lemma 1.3].
We are now in a position to provide a new proof of [RaToVi25, Thm. 0.3] as follows.
Corollary 3.4.
Let ${\mathcal{X}}\hookrightarrow{\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$ be a continuous inclusion of a nonzero complex Hilbert space into the space of tempered distributions on ${\mathbb{R}}^{d}$, satisfying the following condition: there exists a constant $C_{0}>0$ such that for all $f\in{\mathcal{X}}$ and $x,\xi\in{\mathbb{R}}^{d}$ we have $f(\cdot-x){\rm e}^{{\rm i}\langle\cdot,\xi\rangle}\in{\mathcal{X}}$ and $\|f(\cdot-x){\rm e}^{{\rm i}\langle\cdot,\xi\rangle}\|_{\mathcal{X}}\leq C_{0}\|f\|_{\mathcal{X}}$.
Then ${\mathcal{X}}=L^{2}({\mathbb{R}}^{d})$ as vector spaces, and the norms $\|\cdot\|_{\mathcal{X}}$ and $\|\cdot\|_{L^{2}({\mathbb{R}}^{d})}$ are equivalent.
Proof.
We consider the Heisenberg group $G=H_{d}={\mathbb{R}}^{d}\times{\mathbb{R}}^{d}\times{\mathbb{R}}$ with its group operation
$(x,\xi,t)\cdot(y,\eta,s):=(x+y,\xi+\eta,t+s+(\langle\xi,y\rangle-\langle\eta,x\rangle)/2)$
and the Schrödinger representation $\pi\colon H_{d}\to{\mathcal{B}}(L^{2}({\mathbb{R}}^{d}))$ given by
$(\pi(x,\xi,t)f)(y)={\rm e}^{{\rm i}(\langle\xi,y\rangle+t+\langle\xi,x\rangle/2)}f(y+x).$
(See e.g., [CG90, Ex. 2.2.6].) Then we have ${\mathcal{H}}^{\infty}={\mathcal{S}}({\mathbb{R}}^{d})$ hence ${\mathcal{H}}^{-\infty}={\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$. Thus the assumption on $\mathcal{X}$} implies that the hypotheses (H1) and (H2) in Theorem 3.1, as well as (H4) in Remark 3.3 are satisfied. Now the assertion follows by Theorem 3.1 along with Remark 3.3. ∎
4. A result for nilpotent Lie groups
We now assume that $G$ is a 1-connected nilpotent Lie group and ${\mathcal{S}}(G)$ is the associative Fréchet algebra of Schwartz functions on $G$ with the convolution product, hence ${\mathcal{S}}(G)$ is a *-subalgebra of $L^{1}(G)$. Let $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ be an irreducible (continuous) unitary representation. Using the Banach *-algebra representation $\pi\colon L^{1}(G)\to{\mathcal{B}}({\mathcal{H}})$, we have $\pi({\mathcal{S}}(G)){\mathcal{H}}^{\infty}\subseteq{\mathcal{H}}^{\infty}$, and this gives rise to the algebra representation
$\pi^{\infty}\colon{\mathcal{S}}(G)\to{\mathcal{L}}({\mathcal{H}}^{\infty}).$
Also, the map
$\pi^{-\infty}\colon{\mathcal{S}}(G)\to{\mathcal{L}}({\mathcal{H}}^{-\infty}),\quad\pi^{-\infty}(\varphi):=(\pi(\varphi^{*}))^{\prime}$
is an algebra representation.
The maps in (2.1) intertwine the representations $\pi^{\infty}$, $\pi|_{{\mathcal{S}}(G)}$, $\pi^{-\infty}$ of the associative *-algebra ${\mathcal{S}}(G)$ and $\pi^{-\infty}({\mathcal{S}}(G)){\mathcal{H}}^{-\infty}\subseteq{\mathcal{H}}^{\infty}$. By (2.2) we then have
$\pi^{-\infty}({\mathcal{S}}(G)){\mathcal{H}}^{-\infty}={\mathcal{H}}^{\infty}.$ (4.1)
Lemma 4.1.
If $\mathcal{V}$$\subseteq{\mathcal{H}}^{\infty}$ is a $\mathbb{C}$}-linear subspace satisfying $\pi^{\infty}({\mathcal{S}}(G)){\mathcal{V}}\subseteq{\mathcal{V}}$ then either ${\mathcal{V}}=\{0\}$ or $\mathcal{V}$$\mathcal{H}$$^{\infty}$, that is, the representation $\pi^{\infty}\colon{\mathcal{S}}(G)\to{\mathcal{L}}({\mathcal{H}}^{\infty})$ is algebraically irreducible.
Proof.
See [Ho77, Cor. 3.4.1], and also [dC87, Cor. 3.4] and [BeBeGa18, Thm. 2.9]. ∎
Before continuing the preparations for the proof of Theorem 4.4, we derive the following simple result that has a similar flavour and may hold an independent interest.
Proposition 4.2.
In the conditions above, if $\mathcal{Y}$$\subseteq{\mathcal{H}}^{-\infty}$ is a $w^{*}$-closed linear subspace with $\pi^{-\infty}({\mathcal{S}}(G)){\mathcal{Y}}\subseteq{\mathcal{Y}}$, then we have either ${\mathcal{Y}}=\{0\}$ or $\mathcal{Y}$$\mathcal{H}$$^{-\infty}$.
Proof.
Let ${\mathcal{Y}}^{\perp}:=\{v\in{\mathcal{H}}^{\infty}\mid\langle{\mathcal{Y}},v\rangle=\{0\}\}$, which is a closed linear subspace of $\mathcal{H}$$^{\infty}$. Since $\mathcal{Y}$$\subseteq{\mathcal{H}}^{-\infty}$ is a $w^{*}$-closed linear subspace, we obtain a topological linear isomorphism
${\mathcal{Y}}\simeq({\mathcal{H}}^{\infty}/{\mathcal{Y}}^{\perp})^{\prime}.$ (4.2)
On the other hand, the hypothesis $\pi^{-\infty}({\mathcal{S}}(G)){\mathcal{Y}}\subseteq{\mathcal{Y}}$ implies $\pi^{\infty}({\mathcal{S}}(G)){\mathcal{Y}}^{\perp}\subseteq{\mathcal{Y}}^{\perp}$ hence, by Lemma 4.1, we have either ${\mathcal{Y}}^{\perp}=\{0\}$ or $\mathcal{Y}$$^{\perp}$$\mathcal{H}$$^{\infty}$. Now the assertion follows by (4.2). ∎
The next lemma is a more precise form of [Ho77, Cor. 3.4.2]. We give below its short proof. For this statement we recall that ${\mathcal{S}}(G)$ is a *-subalgebra of the Banach algebra $L^{1}(G)$, and by *-representation of ${\mathcal{S}}(G)$ on a Hilbert space $\mathcal{X}$} we mean any *-morphism $\sigma\colon{\mathcal{S}}(G)\to{\mathcal{B}}({\mathcal{X}})$ without any assumptions on continuity.
Lemma 4.3.
Let $\pi$ be as above and $\sigma\colon{\mathcal{S}}(G)\to{\mathcal{B}}({\mathcal{X}})$ be a *-representation on a Hilbert space $\mathcal{X}$}. Assume that the linear map $T\colon{\mathcal{H}}^{\infty}\to{\mathcal{X}}$ satisfies
$T\pi(\varphi)=\sigma(\varphi)T\quad\text{for all }\varphi\in{\mathcal{S}}(G).$
Then $T$ extends to a bounded operator on $\mathcal{H}$} that is a scalar multiple of an isometry.
Proof.
Let $v_{1},v_{2}\in{\mathcal{H}}^{\infty}\setminus\{0\}$ be such that $\langle v_{1},v_{2}\rangle=0$. The orthogonal projection $P:=\langle\cdot,v_{1}\rangle{v_{1}}$ belongs to ${\mathcal{B}}({\mathcal{H}})^{\infty}$ by [BeBe10, Ex. 3.1]. By [Ho77, Thm. 3.4], there is $\varphi\in{\mathcal{S}}(G)$ with $P=\pi(\varphi)$, hence $P=P^{*}=P^{2}=\pi(\psi)$ where $\psi:=\varphi^{*}\ast\varphi\in{\mathcal{S}}(G)$. Then, by hypothesis, $\sigma(\psi)T=TP=\langle\cdot,v_{1}\rangle Tv_{1}$, hence
$\sigma(\psi)Tv_{1}=Tv_{1}\;\;\text{and}\;\;\sigma(\psi)Tv_{2}=0.$
It follows that $Tv_{1}$, respectively $Tv_{2}$, are eigenvectors for the eigenvalues $1$, respectively 0, of the operator $\sigma(\psi)\in{\mathcal{B}}({\mathcal{X}})$. Here $\sigma(\psi)=\sigma(\varphi)^{*}\sigma(\varphi)$ since the mapping $\sigma\colon{\mathcal{S}}(G)\to{\mathcal{B}}({\mathcal{X}})$ is a *-representation, hence $\sigma(\psi)$ is a self-adjoint operator, and then its eigenvectors $Tv_{1}$ and $Tv_{2}$ are orthogonal.
Therefore we have obtained that for $v_{1},v_{2}\in{\mathcal{H}}^{\infty}$,
$\langle v_{1},v_{2}\rangle=0\;\Longrightarrow\;\langle Tv_{1},Tv_{2}\rangle=0.$ (4.3)
Let now $v,w\in{\mathcal{H}}^{\infty}$ be arbitrary with $\|v\|=\|w\|=1$. Set $v_{2}=w-\langle w,v\rangle v$. Then $\langle v,v_{2}\rangle=0$ hence, by (4.3), $\langle Tv,Tv_{2}\rangle=0$ , that is,
$\langle Tv,Tw\rangle=\langle v,w\rangle\|Tv\|^{2}.$ (4.4)
Interchanging the roles of $w$ and $v$, we obtain $\langle v,w\rangle\|Tv\|^{2}=\langle v,w\rangle\|Tw\|^{2}$. It follows that
$\|Tv\|=\|Tw\|\quad\text{if }\langle v,w\rangle\neq 0.$
If $\langle v,w\rangle=0$ and we put $w_{\epsilon}=(w+\epsilon v)/\|w+\epsilon v\|$ for every $\epsilon>0$, then $\langle v,w_{\epsilon}\rangle\neq 0$. Using the above equality we get that
$\|Tv\|=\|w+\epsilon v\|^{-1}\|T(w+\epsilon v)\|=\|w+\epsilon v\|^{-1}\|Tw+\epsilon Tv\|.$
Letting $\epsilon\to 0$, we get that
$\|Tv\|=\|Tw\|\quad\text{for every}\quad v,w\in{\mathcal{H}}^{\infty},\;\|v\|=\|w\|=1.$
Thus $\lambda_{0}:=\|Tv\|$ is independent on $v\in{\mathcal{H}}^{\infty}$ with $\|v\|=1$. Then by (4.4) we get
$\langle Tv,Tw\rangle=\lambda_{0}^{2}\langle v,w\rangle$ (4.5)
for every $v,w\in{\mathcal{H}}^{\infty}$. Since $\mathcal{H}$$^{\infty}$ is dense in $\mathcal{H}$}, (4.5) shows that $T$ extends as a bounded operator on $\mathcal{H}$}, (4.5) holds for $v,w\in{\mathcal{H}}$, hence the extension of $T$ is a scalar multiple of an isometry. ∎
Theorem 4.4.
Let $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ be an irreducible unitary representation of a 1-connected, nilpotent Lie group $G$. Let $\mathcal{X}$$\subseteq{\mathcal{H}}^{-\infty}$ be a $\mathbb{C}$}-linear subspace endowed with the structure of a complex Hilbert space satisfying the hypothesis (H1) and, additionally,
  1. (H5)
    $\pi^{-\infty}({\mathcal{S}}(G)){\mathcal{X}}\subseteq{\mathcal{X}}$.
Then either ${\mathcal{X}}=\{0\}$ or $\mathcal{X}$$\mathcal{H}$}.
Proof.
We start by the preliminary remark that for any unitary representation $\pi$ of an arbitrary Lie group $G$, if $0\neq x\in{\mathcal{H}}^{-\infty}$, then there exists $v\in{\mathcal{H}}^{\infty}$ with $\langle x,v\rangle\neq 0$ hence for a suitable $\varphi\in C_{0}^{\infty}(G)$ we have $\langle x,\pi(\varphi)v\rangle\neq 0$, which implies $\pi^{-\infty}(C_{0}^{\infty}(G))x\neq\{0\}$.
We now come back to the proof and assume ${\mathcal{X}}\neq\{0\}$. Then, by the above remark along with (4.1) and (H5), we have
$\{0\}\neq\pi^{-\infty}({\mathcal{S}}(G)){\mathcal{X}}\subseteq{\mathcal{X}}\cap{\mathcal{H}}^{\infty}.$
On the other hand, the $\mathbb{C}$}-linear space $\pi^{-\infty}({\mathcal{S}}(G)){\mathcal{X}}$ is invariant to $\pi^{-\infty}({\mathcal{S}}(G))$ since $\pi^{-\infty}\colon{\mathcal{S}}(G)\to{\mathcal{L}}({\mathcal{H}}^{-\infty})$ is an algebra representation. Since the inclusion $\mathcal{H}$$^{\infty}\hookrightarrow{\mathcal{H}}^{-\infty}$ intertwines the representations $\pi^{\infty}$ and $\pi^{-\infty}$ of the algebra ${\mathcal{S}}(G)$, an application of Lemma 4.1 then shows that $\pi^{-\infty}({\mathcal{S}}(G)){\mathcal{X}}={\mathcal{H}}^{\infty}$, which further implies ${\mathcal{H}}^{\infty}\subseteq{\mathcal{X}}$. Since the representation $\pi$ of the nilpotent Lie group $G$ is irreducible, it follows by Lemma 4.3 along with hypothesis (H5) that ${\mathcal{H}}\subseteq{\mathcal{X}}$.
However, as in the proof of Theorem 3.1, we have the continuous linear operator $\iota\colon{\mathcal{H}}^{\infty}\to{\mathcal{X}}$ given by
$(\forall v\in{\mathcal{H}}^{\infty},x\in{\mathcal{X}})\quad\langle x,v\rangle=\langle x,\iota(v)\rangle_{\mathcal{X}}.$
with $\iota({\mathcal{H}}^{\infty})$ dense in $\mathcal{X}$}. Since $\iota(v)=v$ for $v\in{\mathcal{H}}^{\infty}\cap{\mathcal{X}}$ and we established above that ${\mathcal{H}}\subseteq{\mathcal{X}}$, we actually obtain ${\mathcal{H}}={\mathcal{X}}$. ∎
Acknowledgment
We wish to thank the Referee for the remarks on our paper and for pointing out [Ne00, Thm. 2.8], and indicating how to considerably improve the paper, by extending of Theorem 3.1 to general Lie groups, as well as further developments along the lines of Theorem 4.4.
We also thank Dr. Shubham Bais for kindly providing an electronic copy of [BaMoPi24].